# LLM-Powered Approximate Intermittent Computing

### v1

@@ 2.2. Dynamic tuning via adjustable knobs
To automate the process of identifying the optimal knob values, we employed Bayesian optimization, which determines knob configurations that achieve the desired balance between energy efficiency and computational accuracy, accounting for the interactions among multiple knobs.
@@ 2.3. Context-aware Chain-of-Thought
Approach: To address these issues, multiple strategies are available, including chain-of-thought (CoT) prompting with few-shot learning, fine-tuning models with additional data, or implementing retrieval-augmented generation (RAG) to contextualize LLM outputs. We dismissed the latter two options due to their reliance on comprehensive datasets that pair original code with approximated versions, which are not readily available. Instead, we adopt a context-aware CoT strategy with few-shot learning to enhance the quality of the LLM’s code output.
@@ 2.4. Workflow
Each design choice in developing CheckMate stemmed from addressing practical challenges and was iteratively refined through experimentation. The integration of feedback loops, the deployment of knobs, the use of Bayesian optimization, and the adoption of a CoT prompting strategy collectively enhance CheckMate’s ability to generate effective approximations for batteryless IoT applications through a one-click solution.
@@ 3.1. User Input
CheckMate requires six essential inputs in JSON format, which users may modify: original application source-code, used to identify approximation opportunities; input traces, representing potential application inputs and energy harvesting traces, or, if unavailable, detailed input characteristics to simulate a trace; accuracy class ($a_{c}$), used to calculate output deviations (cf. Section 3.3 and Table1); and error bound ($e_{b}$), which defines acceptable deviation limits in the output. Additionally, knowledge of the platform and its architecture (e.g., Arm Cortex M or MSP430) aids in precise energy consumption estimation, while the capacitor size determines the minimum energy required to prevent non-progressive states, where the program fails to advance beyond a certain code segment (Hester and Sorber, 2017).
@@ 4. Evaluation
These benchmarks encompass a wide range of programming structures and varying code complexities, enabling a thorough evaluation of CheckMate’s versatility and effectiveness across a spectrum of IoT applications. For instance, the Sobel filter involves computationally intensive image processing, while Activity Recognition employs classification with limited adjustable parameters. The FFT focuses on mathematical computations optimized for low-power microcontrollers, whereas the link estimator handles lightweight network statistics. The Boyer-Moore String Search addresses complex string matching with specific data dependencies, and Bitcount involves bit manipulation with simple control structures.
@@ 4.2. Results → Performance
We evaluated the accuracy of the estimator using percentage absolute error. Given its small and straightforward codebase, this application offered a single approximation opportunity: loop perforation on link’s running history of packet reception rates. As shown in Figure 8(e), CheckMate achieved a 38% reduction in power cycles with a 21% error EWMA. Like the Sobel filter, the fine-tuning in Figure 6 displayed two clusters: one with errors beyond the acceptable threshold and the other within acceptable bounds. Unlike Sobel and other applications, however, this application exhibited a narrower range of power cycles, likely due to its simple computations.
@@ 4.3. Results → Baseline Comparison
We tested CheckMate on two embedded applications from the ACCEPT repository, using the speedup metric—defined as the ratio of clock-cycles in the original to approximated code—to measure processing time reduction, ensuring comparability with ACCEPT’s results.
@@ 4.5. Small-scale user study
We conducted a user study involving seventeen computer science students [^4] This study was approved by our Institutional Review Board. to evaluate the user experience of CheckMate. Our key findings reveal that CheckMate offers two major benefits: 1) it significantly reduces the time required for performing approximations, and 2) it reduces the chances of higher errors in the output compared to manual methods.

### v2

# CheckMate: LLM-Powered Approximate Intermittent Computing

@@ 2.2. Dynamic tuning via adjustable knobs
To automate the process of identifying the optimal knob values, we employed Bayesian optimization, which determines knob configurations that achieve the desired balance between energy efficiency and computational accuracy, accounting for the interactions among multiple knobs. The knobs also serve as mechanisms to mitigate inaccurate approximations by enabling the optimization algorithm to ‘tune out’ [^3] The optimization algorithm can adjust the knob to a value that ensures the corresponding code section behaves as it would in the original, unmodified version, effectively nullifying the approximation. the approximation.
@@ 2.3. Context-aware Chain-of-Thought
Approach: To address these issues, multiple strategies are available, including chain-of-thought (CoT) prompting with few-shot learning, fine-tuning models with additional data, or implementing retrieval-augmented generation (RAG) to contextualize LLM outputs. Due to the absence of sufficiently large and relevant datasets on approximate computing, running a specialized fine-tuned model or one equipped with retrieval-based generation is impractical. Instead, we adopt a context-aware CoT strategy with few-shot learning to enhance the quality of the LLM’s code output. This approach allows us to supplement the LLM’s inherent reasoning with explicit problem-solving steps that align with expert decision-making, guiding the model toward more accurate and contextually appropriate outputs. Additionally, incorporating a few carefully selected examples of approximations helps the LLM recognize how knobs and approximation techniques should be structured within the code
@@ 2.4. Workflow
Each design choice in developing CheckMate stemmed from addressing practical challenges and was iteratively refined through experimentation. The integration of feedback loops, the deployment of knobs, the use of Bayesian optimization, and the adoption of a CoT prompting strategy collectively enhance CheckMate’s ability to generate effective approximations while minimizing errors, for batteryless IoT applications through a one-click solution.
@@ 3.1. User Input
CheckMate requires six essential inputs in JSON format, which users may modify: original application source-code, used to identify approximation opportunities; input traces, representing potential application inputs and energy harvesting traces, or, if unavailable, detailed input characteristics to simulate a trace; accuracy class ($a_{c}$), used to calculate output deviations (cf. Section 3.3 and Table1); and error bound ($e_{b}$), which defines acceptable deviation limits in the output. Additionally, knowledge of the platform and its architecture (e.g., Arm Cortex M or MSP430) aids in precise energy consumption estimation, while the capacitor size determines the minimum energy required to prevent non-progressive states, where the program fails to advance beyond a certain code segment (Hester and Sorber, 2017). An optional parameter, LLM temperature, can be adjusted, with a default low value (0–0.3), as this range minimizes the LLM’s randomness and is generally preferred for coding tasks (Chen et al., 2021).
@@ 4. Evaluation
Table 2. Application size. The reported sizes exclude storage used by the checkpointing strategy, which can add up to 58 kB of Flash. Token counts represent the number of tokens required to tokenize the codebase.
Application Total Flash Usage Total RAM Usage Token Count Line of Code
Sobel 9.0 kB 3712 B 1002 123
Activity Recognition 1200 B 274 B 2213 307
FFT 11413 B 26 B 911 119
String Search 4324 B 942 B 689 94
Link Estimator 9600 B 2516 B 403 78
Bitcount 574 B 10 B 489 80
As shown in Table 2, these benchmarks encompass a wide range of programming structures and varying code complexities, enabling a thorough evaluation of CheckMate’s versatility and effectiveness across a spectrum of IoT applications. For instance, the Sobel filter involves computationally intensive image processing, while Activity Recognition employs classification with limited adjustable parameters. The FFT focuses on mathematical computations optimized for low-power microcontrollers, whereas the link estimator handles lightweight network statistics. The Boyer-Moore String Search addresses complex string matching with specific data dependencies, and Bitcount involves bit manipulation with simple control structures.
@@ 4.2. Results → Performance
We evaluated the accuracy of the estimator using percentage absolute error. Given its straightforward codebase, this application offered a single approximation opportunity: loop perforation on link’s running history of packet reception rates. As shown in Figure 8(e), CheckMate achieved a 38% reduction in power cycles with a 21% error EWMA. Like the Sobel filter, the fine-tuning in Figure 6 displayed two clusters: one with errors beyond the acceptable threshold and the other within acceptable bounds. Unlike Sobel and other applications, however, this application exhibited a narrower range of power cycles, likely due to its simple computations.
@@ 4.3. Results → Baseline Comparison
We tested CheckMate on the two embedded applications, Sobel and Activity Recognition, from the ACCEPT repository, using the speedup metric—defined as the ratio of clock cycles in the original to approximated code—to measure processing time reduction, ensuring comparability with ACCEPT’s results.
We selected these two applications to ensure a direct and fair comparison with ACCEPT. The developers of ACCEPT evaluated and reported results exclusively on these two embedded system applications, making them the only available baseline for comparison. Expanding our benchmark set beyond these applications would have introduced methodological inconsistencies, reducing the validity of direct performance comparisons.
@@ 4.5. Small-scale user study
We conducted a user study involving seventeen computer science students to evaluate the user experience of CheckMate. Our key findings reveal that CheckMate offers two major benefits: 1) it significantly reduces the time required for performing approximations, and 2) it reduces the chances of higher errors in the output compared to manual methods.
@@
5. Discussion and Limitations
The presented solution approach also introduces certain limitations and challenges related to the performance and practicability of CheckMate. In this section, we highlight some of these shortcomings and discuss potential solutions that users may adopt.
Approximation overhead: Several approximations introduce computational overhead. For example, a loop perforation approximation may add extra floating-point arithmetic operations, which contribute to the overall computation of the application. On its own, this overhead may not significantly alter the number of checkpoints required for the application to run to completion. However, when a large number of knobs are created by the LLM, it is possible for these knobs to collectively generate a significant overhead.
This overhead becomes particularly significant when many of the created knobs are ‘tuned out’ by the Bayesian optimizer. A knob is considered tuned out when its setting ensures that the corresponding section of code behaves identically to the original application code. Consequently, tuned-out knobs introduce unnecessary computational overhead.
Much of this issue is mitigated by compilers, which automatically eliminate redundant code. However, there are instances where the compiler fails to remove such overhead. These typically arise when the LLM provides knob ranges 5 that lack a value capable of making the program behave exactly as the original code. In such cases, the Bayesian optimizer selects the knob value closest to ”zero” (i.e., the setting most similar to the original program operation), but the compiler remains unable to eliminate the associated overhead.
Privacy concerns: Using commercial APIs for testing and evaluation of CheckMate can raise privacy concerns for developers operating in proprietary or confidential environments. To address this, CheckMate is designed to seamlessly integrate with any LLM API via LangChain, enabling it to operate with locally deployed LLMs. With the rapid advancements in open-source LLMs, such as DeepSeek’s (DeepSeek-AI, 2024), many of which have reached performance levels comparable to commercial alternatives while requiring less computational power and budget, it is reasonable to assert that CheckMate would deliver similar performance when leveraging these open-source models.
LLM Specifications and Application Size Dynamics: The relationship between application size and the number of model parameters is a critical determinant of performance. As shown in Figure 5, CheckMate’s effectiveness declines significantly when using smaller LLMs. For instance, while both GPT-4o and GPT-4o Mini share the same context window size, they differ in parameter count. The smaller GPT-4o Mini struggles with larger and more complex applications, exhibiting noticeable performance degradation. Additionally, it fails to capture some of the optimal approximations identified by its larger counterpart.
Another key factor is the relationship between application size and context window length. Larger applications require a longer context window for effective processing. To ensure broad generalizability across various LLMs while maintaining efficiency, CheckMate was designed to fit within a 16k-token context window [^7] Different LLMs use varying tokenization strategies..